Notational Inc. is considering installing a new server. The machine costs $100,000 and is expected to have a useful economic life of 8 years, after which it will have a book value of $0. In addition to the equipment costs, management expects installation costs of $10,000 and an initial outlay for net working capital of $12,000. The new server is expected to generate an additional $10,000 per year in earnings after tax over its useful life, but an additional $5,000 per year is required in net working capital. Net working capital will be recovered at the end of 8 years. Assume that National has a cost of capital of 10%. Calculate the NPV and IRR for this project.
Answer #1

Cash Flow= 10000-5000+12500 (Depreciation)=17500

At the end of the 8th year Wc is returned=12000+5000*8=52000

Initial Investment= 100000+10000+12000=122000

Current value of CF= 17500*((1/1.1)+....+(1/1.1^8))+(52000/1.1^8)=17500*11.81+24299=230974

NPV=230974-122000=108974